# Quadratic counts of highly tangent lines to hypersurfaces

### v1

@@ Abstract.
We give two geometric interpretations for the local type of a line that is highly tangent to a hypersurface in a single point. One interpretation is phrased in terms of the Wronski map, while the other interpretation relates to the fundamental forms of the hypersurface. These local types are the local contributions of an quadratic form-valued Euler number that depends on a choice of orientation.
@@ 1. Introduction
For polynomials of even degree, this signed count of real roots is always 0. For odd polynomials, the signed count is ±$\pm 1$. This sign ambiguity can be interpreted as an artifact of non-orientability: the roots of a polynomial correspond to the intersection of two plane curves, and Bézout’s theorem is not an orientable problem over non-closed fields [McK21].
By highly tangent line we mean a line meeting a hypersurface in $\mathbb{P}^{n}$ at a point with contact order $12n-$ 1. In this article, we will investigate the geometric weight with which to count highly tangent lines to the hypersurface. Analogous to the case of real roots of odd polynomials, this enumerative problem is not orientable in many cases of interest, so the total count of highly tangent lines depends on the choice of hypersurface. Despite the lack of invariant total count, we find that this counting weight admits two interesting descriptions.
@@
Let $k$ be a field. Let $X=\mathbb{V}(F)$ by a smooth, general hypersurface in $\mathbb{P}^{n}_{k}$ of degree $d\geq 2n-1$. Let $\Phi$ be the flag variety of pointed lines in $\mathbb{P}^{n}_{k}$, and let $\beta\colon\Phi\to\mathbb{P}^{n}_{k}$ send a pointed line to its marked point.
Let $L$ be a line such that the intersection multiplicity of $L$ and $X$ at $p$ is $12n-$ 1. Then the local index associated to $L$ is given by the Wronskian determinant (with respect to the local parameter of $L$ at $p$) of the gradient $\nabla\beta^{*}F$.
@@ 1.1. Outline
The enumerative problem of counting highly tangent (projective) linear spaces to hypersurfaces can be formulated in terms of a bundle of principal parts over a flag variety of pointed linear spaces (see e.g. [EH16, §11.1]). In Section 2, we will recall the construction of this flag variety, which serves as the parameter space of objects that we will count. We will also discuss standard coordinates on the flag variety. In Section 3, we will recall the bundle of principal parts, as well as convenient local trivializations of this bundle. We will also make a few remarks about Euler numbers in this section.
@@ Acknowledgements
We thank Ethan Cotterill, Steven Kleiman, Marc Levine, Wenbo Niu, Sabrina Pauli, and Felipe Voloch for useful discussions. SM received support from an NSF MSPRF grant (DMS-2202825). GM is a member of GNSAGA (INdAM), and is supported by FCT - Fundação para a Ciência e a Tecnologia, under the project: UIDP/04561/2020 (https://doi.org/10.54499/UIDP/04561/2020). WO thanks DJ Merril for help with implementing several of the computations in [MMO24].
@@ 2. The parameter space
Since $\Phi$ and $G$ are both smooth, we have
@@
By the relative Euler sequence
$0\to\mathrm{T}_{\Phi/G}\to\mathrm{T}_{\Phi}\to\pi^{*}\mathrm{T}_{G}\to 0,$
we have $\det\mathrm{T}_{\Phi/G}\cong\det\mathrm{T}_{\Phi}\otimes\pi^{*}\det\mathrm{T}^{\vee}_{G}$. Taking duals and applying Lemma 2.4, we find
@@ 2.1. Coordinates
More precisely, let $11\leq i_{1}<\ldots<i_{r+1}\leq n+$ 1, and denote $I=\{i_{1},\ldots,i_{r+1}\}$. Let $\{c_{1},\ldots,c_{n-r}\}=\{1,\ldots,n+1\}-I$, with $c_{1}<\ldots<c_{n-r}$. Let $U_{I}\subset\mathbb{G}(r,n)$ denote the open set of $(r+1)$-planes such that $\det((\tilde{x}_{t,i_{s}})_{t,s=1}^{r+1})\neq 0$ (in the parameterization given in Equation 2.1). By row reducing, all such $(r+1)$-planes in $U_{I}$ can be obtained by taking $(\tilde{x}_{t,i_{s}})_{t,s=1}^{r+1}$ to be the identity matrix. We thus obtain an isomorphism
@@
Let $U_{I,\ell}\subset\mathbb{P}\mathcal{S}$ be the open set of pairs $(H,\mathrm{span}(v))$, where $H\in U_{I}$ and the coordinate of $\tilde{e}_{i_{\ell}}$ in $v$ (from Equation 2.2) is non-zero. Let $\varphi_{I,\ell}:=\varphi_{I}\times\psi_{\ell}$, where $\psi_{\ell}:\{[y_{1}:\cdots:y_{r+1}]:y_{\ell}\neq 0\}\to\mathbb{A}^{r}$ are the twisted affine charts on $\mathbb{P}^{r}$ defined in [McK21, p. 638]. The standard coordinates for $\Phi_{r,n}\cong\mathbb{P}\mathcal{S}$ consist of the local coordinates $\{(U_{I,\ell},\varphi_{I,\ell})\}_{I,\ell}$.
@@ 3. The bundle
We will briefly revise the definition of the flag variety $\Phi_{r,n}$ parameterizes projective $r$-planes in $\mathbb{P}^{n}$ equipped with a point. We now need a vector bundle encoding the condition that such a pointed plane $(H,p)$ meets a hypersurface at the point $p$ to a prescribed order. Such a bundle is known as the bundle of principal parts [Gro67, Définition (16.3.1)].
@@
(3.1) $0\longrightarrow\mathcal{O}_{\Phi}(d)\otimes\mathrm{Sym}^{m-1}(\Omega_{\Phi/G})\longrightarrow\mathcal{E}_{m}\longrightarrow\mathcal{E}_{m-1}\longrightarrow 0$
It follows (by induction) that $\operatorname{rank}\mathcal{E}_{m}=\binom{r+m-1}{r}$ (see e.g. [Vai81, (3.5) Proposition]). We can now use Equation 3.1 to compute $\det\mathcal{E}_{m}$.
@@
using standard computations (see e.g. [Har77, Exercise II.5.16(c)] and [Ful98, B.5.8]). The desired result now follows from the addition rules of binomial coefficients. ∎
@@
The Picard group of $\Phi$ is generated by $\pi^{*}\mathcal{O}_{G}(1)$ and $\mathcal{O}_{\Phi}(1)$ (see e.g. [Har77, Exercise II.7.9(a)]), so $\mathcal{E}_{m}$ is relatively orientable if and only if
@@
One can check that $\mathcal{E}_{m}$ is not orientable if $r=1$, as pointed out in [Mur24a, Remark 6.4]. Using the code of [MMO24], we found that for $r,m,n\leq 5000$ and $r<n$, the conditions of Proposition 3.4 are only satisfied by tuples $(d,r,m,n)$ with $d$ even and $(r,m,n)$ one of the following:
@@
Euler numbers of vector bundles over odd dimensional schemes are hyperbolic [SW21, Proposition 19], so $e(\mathcal{E}_{m})$ is some multiple of $\langle 1\rangle+\langle-1\rangle$ whenever $\dim\Phi$ is odd. Since $\operatorname{rank}e(\mathcal{E}_{m})$ is equal to the complex count (i.e. the top Chern number of $\mathcal{E}_{m}$), the desired formula will follow if we can show that $\dim\Phi$ is odd in the relatively orientable case. We will show the contrapositive.
Let $B=\binom{r+m-1}{m-2}$. If $\dim\Phi=r(n-r)+n$ is even, then $n$ and $r$ must both be even. Thus $d(r(n-r)+n)$ must be even, so $(r+1)(B-1)$ must be even as well. It follows that $B$ must be odd. But this contradicts $B\equiv n\mod 2$, so $\mathcal{E}_{m}\to\Phi$ cannot be relatively orientable. ∎
@@ 3.1. Orienting divisors
Let $X$ be a a smooth proper $k$-scheme $X\to\operatorname{Spec}k$. A vector bundle $V$ over $X$ is said to be relatively orientable relative to a divisor $D\subset X$ if there exists a line bundle $\mathcal{L}$ and an isomorphism $\rho\colon\det V\otimes\omega_{X/k}\otimes\mathcal{O}(D)\to\mathcal{L}^{\otimes 2}$. We will refer to such a divisor as an orienting divisor for $V\to X$.
To find a suitable orienting divisor for $\mathcal{E}_{m}\to\Phi$, we focus on our case of interest. That is, set $r=1$ and assume that $\operatorname{rank}\mathcal{E}_{m}=\dim\Phi$. This gives us an equality $2n-1=\binom{m}{m-1}$, so we find that $m=2n-1$. Under these assumptions, Lemma 3.3 implies that
@@ 3.2. Trivializations
Our preferred coordinates on $\Phi$ give us a natural choice of trivialization for $\mathcal{O}_{\Phi}(D)$. The proof of the following lemma is standard (c.f. [Mur24a, Lemma 3.8 and Proposition 5.1]), so we opt to omit it.
@@
Since $r=1$, we have that $\mathcal{O}_{\Phi}(d)\otimes\mathrm{Sym}^{a}(\Omega_{\Phi/G})$ is a line bundle for $10\leq a\leq m-$ 1. We can obtain a local trivialization
@@
by locally trivializing the line bundles $\mathcal{O}_{\Phi}(d)\otimes\mathrm{Sym}^{a}(\Omega_{\Phi/G})$ for $10\leq a\leq m-$ 1. Recall that for any finite rank vector bundle $V$, we have a canonical isomorphism
@@
[…]
@@
To conclude, let $m=2n-1$ (so that $\operatorname{rank}\mathcal{E}_{m}=\dim\Phi_{1,n}$). We need to show that our local coordinates and local trivializations are compatible with our relative orientation of $\mathcal{E}_{2n-1}$ relative to $D$. This will immediately follow from Lemma 3.9 once we prove the following lemma.
@@
Over the field of complex numbers, the Euler number of $\mathcal{E}_{m}\to\Phi$ is well-known. See for example [BW25], and [Mur24b, Equation (3.4)] for an approach using Gromov–Witten invariants.
@@ 4. Wronskian interpretation of the local index
In this section, we give a more precise statement of Theorem 1.1, which we then prove. Throughout, we assume $r=1$ and $m=2n-1$ (as in Section 3.1). Let $F$ be a global section of $\mathcal{O}_{\mathbb{P}^{n}}(d)$. Note that $\beta^{*}\mathcal{O}_{\mathbb{P}^{n}}(d)\cong\mathcal{O}_{\Phi}(d)$, where $\beta\colon\Phi\to\mathbb{P}^{n}$ is the projection map arising from Definition 2.1. The degree $d$ hypersurface $\mathbb{V}(F)\subset\mathbb{P}^{n}$ determines a section
@@
Let $F$ be a global section of $\mathcal{O}_{\mathbb{P}^{n}}(d)$ with $d\geq 2n-1$. Let $L$ be a line that is tangent to $\mathbb{V}(F)$ at $p\in L$ to order $12n-$ 1. Finally, assume that $(L,p)$ is geometrically reduced as a point in $\Phi$, and that $k(L,p)/k$ is separable. Then
@@
By [BBM+21], we can assume that $k(L,p)=k$ by base changing and tracing back down (if necessary). If $(L,p)\in U_{I,\ell}$, then we have
@@
for $11\leq i\leq n-$ 1. Note that $y$ is precisely the dehomogenization of our local parameter $t$, so that $D^{(a)}_{t}=D^{(a)}_{y}$ on $U_{(n,n+1),2}$.
Let $[z_{1}:\ldots:z_{n+1}]$ denote coordinates on $\mathbb{P}^{n}$. In our coordinates on $U_{(n,n+1),2}$, the map $\beta\colon\Phi\to\mathbb{P}^{n}$ is given by
@@
Brazelton has some results on geometrically interpreting Wronskians in enriched enumerative geometry [Bra25]. The Hasse–Wronskian also arises as the local index for the enriched count of inflection points to superelliptic curves [CDGL+21], showing that our Theorem 1.1 is compatible with this result.
@@ 5. Highly tangent lines and the second fundamental form
(5.1) $0\longrightarrow I^{m}/I^{m+1}\longrightarrow(A\otimes_{k}A)/I^{m+1}\overset{\pi_{m}}{\longrightarrow}(A\otimes_{k}A)/I^{m}\longrightarrow 0$
We call the following map of left $A$-modules the Taylor series map (see [Per95, Proposition 6.4.3])
@@
called $m^{\mathrm{th}}$ fundamental form. We have $\ker(\delta_{0})\cong\Omega_{\mathbb{P}^{n}}|_{X}$ and $\ker(\delta_{1})\cong N^{\vee}_{X/\mathbb{P}^{n}}$ (see [Har77, Theorem II.8.13 and Proposition II.8.12]). If $X$ is a smooth projective variety, those maps naturally glue together on the affine patches of $X$. We are interested in one particular fundamental form.
@@
$\mathcal{F}:=\operatorname{Hom}(N^{\vee}_{X/\mathbb{P}^{n}},\mathrm{Sym}^{2}\Omega_{X})\cong N_{X/\mathbb{P}^{n}}\otimes\mathrm{Sym}^{2}\Omega_{X}.$
@@
The second fundamental form can be constructed by applying [AK70, Section I.3] to the exact sequence (e.g. [Har77, Theorem II.8.17 (2)]):
(5.4) $0\longrightarrow N^{\vee}_{X/\mathbb{P}^{n}}\longrightarrow\Omega_{\mathbb{P}^{n}}|_{X}\longrightarrow\Omega_{X}\longrightarrow 0$
In [EN23, MP24], the $m^{\mathrm{th}}$ fundamental forms are constructed using the Taylor series map applied to the bundles of principal part $\mathcal{P}_{X/k}^{m}(\mathcal{O}_{X})$ [Gro67, Définition (16.3.6)], and using again the properties of the kernel. They suppose $k$ to be algebraically closed, but this hypothesis is not really necessary, at least in order to define the fundamental forms [^2] Confirmed by Wenbo Niu by a personal communication.. Our approach follows [Per95, Appendix A].
@@
In Proposition 5.3, we cannot drop the hypothesis that $X$ admits a theta characteristic. For example, a non-singular real plane curve of degree $4$ with no real points has no theta characteristics [Ati71, Section 5]. See also [GH81, Section 5] for a thorough discussion about theta characteristics of real curves, and [Rön10] for a characterization of the existence of theta characteristic in terms of the splitting of the stable motivic homotopy type of the curve. If $k$ is algebraically closed, then any non-singular curve admits theta characteristics [Mum71, Section 4].
@@
For arbitrary fields note that the functor $\otimes_{k}\bar{k}$ is exact, so we may suppose $k$ algebraically closed. Indeed, Serre’s intersection formula is preserved, see [Har77, Appendix A]. A geometric interpretation of degenerate locus of   II as the locus of inflectional points is given by [MP24, Section 3] (see also [Per95, Appendix B, Theorem 2.3] and [Pie77]).
@@
Since $\mathcal{F}\to X$ is a line bundle over a curve, its Euler number is hyperbolic by [SW21, Proposition 19]. It thus suffices to compute the rank of $e(\mathcal{F})$, which is given by the integral Euler number over an algebraically closed field. The claim now follows from the classical count of $3d(d-2)$ inflectional points to a smooth complex plane curve of degree $d$ [Har77, Exercise IV.2.3]. ∎
@@
The latter equation follows from the first by the Poincaré–Hopf Theorem (see e.g. [BW23, Theorem 1.1]) and the fact that $\langle 3\rangle\mathbb{H}=\mathbb{H}$, so it remains to treat the first equation. Using the same notation at the beginning of this section, assume $X=\operatorname{Spec}(A)$ is localized at a smooth $k$-point. The ideal $I$ is free of rank one, so it is generated by $\mathrm{d}z:=z\otimes 1-1\otimes z$ for some $z\in A$. There exists an isomorphism

### v2

@@ Abstract.
We give two geometric interpretations for the local type of a line that is highly tangent to a hypersurface in a single point. One interpretation is phrased in terms of the Wronski map, while the other interpretation relates to the fundamental forms of the hypersurface. These local types are the local contributions of a quadratic form-valued Euler number that depends on a choice of orientation.
@@ 1. Introduction
For polynomials of even degree, this signed count of real roots is always 0. For polynomials of odd degree, the signed count is $\pm 1$. This sign ambiguity can be interpreted as an artifact of non-orientability: the roots of a polynomial correspond to the intersection of two plane curves, and Bézout’s theorem is not an orientable problem over non-closed fields [MR4211099].
By highly tangent line we mean a line meeting a hypersurface in $\mathbb{P}^{n}$ at a point with contact order at least 2n12n-1. In this article, we will investigate the geometric weight with which to count highly tangent lines to the hypersurface. Analogous to the case of real roots of odd polynomials, this enumerative problem is not orientable in many cases of interest, so the total count of highly tangent lines depends on the choice of hypersurface. Despite the lack of invariant total count, we find that this counting weight admits two interesting descriptions.
@@
Let $k$ be a field. Let $X=\mathbb{V}(F)$ be a smooth, general hypersurface in $\mathbb{P}^{n}_{k}$ of degree $d\geq 2n-1$. Let $\Phi$ be the flag variety of pointed lines in $\mathbb{P}^{n}_{k}$, and let $\beta\colon\Phi\to\mathbb{P}^{n}_{k}$ send a pointed line to its marked point.
Let $L$ be a line such that the intersection multiplicity of $L$ and $X$ at $p$ is 2n12n-1. Then the local index associated to $L$ is given by the Wronskian determinant (with respect to the local parameter of $L$ at $p$) of the gradient $\nabla\beta^{*}F$.
@@ 1.1. Outline
The enumerative problem of counting highly tangent (projective) linear spaces to hypersurfaces can be formulated in terms of a bundle of principal parts over a flag variety of pointed linear spaces (see e.g. [3264, §11.1]). In Section 2, we will recall the construction of this flag variety, which serves as the parameter space of objects that we will count. We will also discuss standard coordinates on the flag variety. In Section 3, we will recall the bundle of principal parts, as well as convenient local trivializations of this bundle. We will also make a few remarks about Euler numbers in this section.
@@
Data management statement
The data that support the findings of this study are openly available in highly_tangent at https://github.com/wgabrielong/highly_tangent/tree/main.
Conflict of interest statement
The authors declare no conflict of interest related to this study.
@@ Acknowledgements
We thank Ethan Cotterill, Steven Kleiman, Marc Levine, Wenbo Niu, Sabrina Pauli, and Felipe Voloch for useful discussions. SM received support from an NSF MSPRF grant (DMS-2202825). GM is a member of GNSAGA (INdAM), and is supported by FCT - Fundação para a Ciência e a Tecnologia, under the project: UID/04561/2025. WO thanks DJ Merril for help with implementing several of the computations in [GabRep].
@@ 2. The parameter space
Since $\Phi$ and $G$ are both smooth, we have $\operatorname{rank}\Omega_{\Phi}=\dim\Phi$ and $\operatorname{rank}\Omega_{G}=\dim{G}$. The rank of the relative cotangent bundle is thus
@@
By the dual relative Euler sequence
$0\to\pi^{*}\Omega_{G}\to\Omega_{\Phi}\to\Omega_{\Phi/G}\to 0,$
we have $\det\Omega_{\Phi/G}\cong\omega_{\Phi}\otimes\pi^{*}\omega_{G}^{\vee}$. Applying Lemma 2.4, we find
@@ 2.1. Coordinates
More precisely, let $1\leq i_{1}<\ldots<i_{r+1}\leq n+1$, and denote $I=\{i_{1},\ldots,i_{r+1}\}$. Let $\{c_{1},\ldots,c_{n-r}\}=\{1,\ldots,n+1\}-I$, with $c_{1}<\ldots<c_{n-r}$. Let $U_{I}\subset\mathbb{G}(r,n)$ denote the open set of $(r+1)$-planes such that $\det((\tilde{x}_{t,i_{s}})_{t,s=1}^{r+1})\neq 0$ (in the parameterization given in Equation 2.1). By row reducing, all such $(r+1)$-planes in $U_{I}$ can be obtained by taking $(\tilde{x}_{t,i_{s}})_{t,s=1}^{r+1}$ to be the identity matrix. We thus obtain an isomorphism
@@
Let $U_{I,\ell}\subset\mathbb{P}\mathcal{S}$ be the open set of pairs $(H,\mathrm{span}(v))$, where $H\in U_{I}$ and the coordinate of $\tilde{e}_{i_{\ell}}$ in $v$ (from Equation 2.2) is non-zero. Let $\varphi_{I,\ell}:=\varphi_{I}\times\psi_{\ell}$, where $\psi_{\ell}:\{[y_{1}:\cdots:y_{r+1}]:y_{\ell}\neq 0\}\to\mathbb{A}^{r}$ are the twisted affine charts on $\mathbb{P}^{r}$ defined in [MR4211099, p. 638]. The standard coordinates for $\Phi_{r,n}\cong\mathbb{P}\mathcal{S}$ consist of the local coordinates $\{(U_{I,\ell},\varphi_{I,\ell})\}_{I,\ell}$.
@@ 3. The bundle
We will briefly revise the definition of the flag variety $\Phi_{r,n}$ parameterizing projective $r$-planes in $\mathbb{P}^{n}$ equipped with a point. We now need a vector bundle encoding the condition that such a pointed plane $(H,p)$ meets a hypersurface at the point $p$ to a prescribed order. Such a bundle is known as the bundle of principal parts [MR238860, Définition (16.3.1)].
@@
(3.1) $0\longrightarrow\mathcal{O}_{\Phi}(d)\otimes\mathrm{Sym}^{m-1}(\Omega_{\Phi/G})\longrightarrow\mathcal{E}_{m}\longrightarrow\mathcal{E}_{m-1}\longrightarrow 0.$
It follows (by induction) that $\operatorname{rank}\mathcal{E}_{m}=\binom{r+m-1}{r}$ (see e.g. [MR0626480, (3.5) Proposition]). We can now use Equation 3.1 to compute $\det\mathcal{E}_{m}$.
@@
using standard computations (see e.g. [MR0463157, Exercise II.5.16(c)] and [MR1644323, B.5.8]). The desired result now follows from the addition rules of binomial coefficients. ∎
@@
The Picard group of $\Phi$ is generated by $\pi^{*}\mathcal{O}_{G}(1)$ and $\mathcal{O}_{\Phi}(1)$ (see e.g. [MR0463157, Exercise II.7.9(a)]), so $\mathcal{E}_{m}$ is relatively orientable if and only if
@@
One can check that $\mathcal{E}_{m}$ is not orientable if $r=1$, as pointed out in [osculating, Remark 6.4]. Using the code of [GabRep], we found that for $r,m,n\leq 5000$ and $r<n$, the conditions of Proposition 3.4 are only satisfied by tuples $(d,r,m,n)$ with $d$ even and $(r,m,n)$ one of the following:
@@
Euler numbers of vector bundles over odd dimensional schemes are hyperbolic [FourLines, Proposition 19], so $e(\mathcal{E}_{m})$ is some multiple of $\langle 1\rangle+\langle-1\rangle$ whenever $\dim\Phi$ is odd. Since $\operatorname{rank}e(\mathcal{E}_{m})$ is equal to the complex count (i.e. the top Chern number of $\mathcal{E}_{m}$), the desired formula will follow if we can show that $\dim\Phi$ is odd in the relatively orientable case. We will show the contrapositive.
Let $B=\binom{r+m-1}{m-2}$. If $\dim\Phi=r(n-r)+n$ is even, then $n$ and $r$ must both be even. Thus d(r(nr)+n)d(r(n-r)+n) must be even, so (r+1)(B1)(r+1)(B-1) must be even as well. It follows that $B$ must be odd. But this contradicts $B\equiv n\mod 2$, so $\mathcal{E}_{m}\to\Phi$ cannot be relatively orientable. ∎
@@ 3.1. Orienting divisors
Let $X\to\operatorname{Spec}{k}$ be a smooth, proper $k$-scheme. A vector bundle $V$ over $X$ is said to be relatively orientable relative to a divisor $D\subset X$ if there exists a line bundle $\mathcal{L}$ and an isomorphism $\rho\colon\det V\otimes\omega_{X/k}\otimes\mathcal{O}(D)\to\mathcal{L}^{\otimes 2}$. We will refer to such a divisor as an orienting divisor for $V\to X$.
To find a suitable orienting divisor for $\mathcal{E}_{m}\to\Phi$, we focus on our case of interest. That is, set $r=1$ and assume that $\operatorname{rank}\mathcal{E}_{m}=\dim\Phi$. This gives us an equality $2n-1=\binom{m}{m-1}$, so we find that m=2n1m=2n-1. Under these assumptions, Lemma 3.3 implies that
@@ 3.2. Trivializations
Our preferred coordinates on $\Phi$ give us a natural choice of trivialization for $\mathcal{O}_{\Phi}(D)$. The proof of the following lemma is standard (c.f. [osculating, Lemma 3.8 and Proposition 5.1]), so we opt to omit it.
@@
Since $r=1$, we have that $\mathcal{O}_{\Phi}(d)\otimes\mathrm{Sym}^{a}(\Omega_{\Phi/G})$ is a line bundle for $0\leq a\leq m-1$. We can obtain a local trivialization
@@
by locally trivializing the line bundles $\mathcal{O}_{\Phi}(d)\otimes\mathrm{Sym}^{a}(\Omega_{\Phi/G})$ for $0\leq a\leq m-1$. Recall that for any finite rank vector bundle $V$, we have a canonical isomorphism
@@
where $\Gamma^{a}$ denotes $a\textsuperscript{th}$ divided powers and $(-)^{\vee}$ denotes the dual. In particular, we have $\mathrm{Sym}^{a}(\Omega_{\Phi/G})\cong\Gamma^{a}(\mathrm{T}_{\Phi/G})^{\vee}$, where $\mathrm{T}_{\Phi/G}$ is the relative tangent bundle. As
@@
To conclude, let m=2n1m=2n-1 (so that $\operatorname{rank}\mathcal{E}_{m}=\dim\Phi_{1,n}$). We need to show that our local coordinates and local trivializations are compatible with our relative orientation of $\mathcal{E}_{2n-1}$ relative to $D$. This will immediately follow from Lemma 3.9 once we prove the following lemma.
@@
Over the field of complex numbers, the Euler number of $\mathcal{E}_{m}\to\Phi$ is well-known. See for example [bertone], and [mu, Equation (3.4)] for an approach using Gromov–Witten invariants.
@@ 4. Wronskian interpretation of the local index
In this section, we give a more precise statement of Theorem 1.1, which we then prove. Throughout, we assume $r=1$ and m=2n1m=2n-1 (as in Section 3.1). Let $F$ be a global section of $\mathcal{O}_{\mathbb{P}^{n}}(d)$. Note that $\beta^{*}\mathcal{O}_{\mathbb{P}^{n}}(d)\cong\mathcal{O}_{\Phi}(d)$, where $\beta\colon\Phi\to\mathbb{P}^{n}$ is the projection map arising from Definition 2.1. The degree $d$ hypersurface $\mathbb{V}(F)\subset\mathbb{P}^{n}$ determines a section
@@
Let $F$ be a global section of $\mathcal{O}_{\mathbb{P}^{n}}(d)$ with $d\geq 2n-1$. Let $L$ be a line that is tangent to $\mathbb{V}(F)$ at $p\in L$ to order 2n12n-1. Finally, assume that $(L,p)$ is geometrically reduced as a point in $\Phi$, and that $k(L,p)/k$ is separable. Then
@@
By [BBMMO21], we can assume that $k(L,p)=k$ by base changing and tracing back down (if necessary). If $(L,p)\in U_{I,\ell}$, then we have
@@
for $1\leq i\leq n-1$. Note that $y$ is precisely the dehomogenization of our local parameter $t$, so that $D^{(a)}_{t}=D^{(a)}_{y}$ on $U_{(n,n+1),2}$.
Let $z_{1}:\ldots:z_{n+1}]$ denote coordinates on $\mathbb{P}^{n}$. In our coordinates on $U_{(n,n+1),2}$, the map $\beta\colon\Phi\to\mathbb{P}^{n}$ is given by
@@
Brazelton has some results on geometrically interpreting Wronskians in enriched enumerative geometry [Bra25]. The Hasse–Wronskian also arises as the local index for the enriched count of inflection points to superelliptic curves [CDGLHS21], showing that our Theorem 1.1 is compatible with this result.
@@ 5. Highly tangent lines and the second fundamental form
(5.1) $0\longrightarrow I^{m}/I^{m+1}\longrightarrow(A\otimes_{k}A)/I^{m+1}\overset{\pi_{m}}{\longrightarrow}(A\otimes_{k}A)/I^{m}\longrightarrow 0.$
We call the following map of left $A$-modules the Taylor series map (see [MR1308020, Proposition 6.4.3])
@@
called $m^{\mathrm{th}}$ fundamental form. We have $\ker(\delta_{0})\cong\Omega_{\mathbb{P}^{n}}|_{X}$ and $\ker(\delta_{1})\cong N^{\vee}_{X/\mathbb{P}^{n}}$ (see [MR0463157, Theorem II.8.13 and Proposition II.8.12]). If $X$ is a smooth projective variety, those maps naturally glue together on the affine patches of $X$. We are interested in one particular fundamental form.
@@
$\mathcal{F}:=\mathcal{H}om(N^{\vee}_{X/\mathbb{P}^{n}},\mathrm{Sym}^{2}\Omega_{X})\cong N_{X/\mathbb{P}^{n}}\otimes\mathrm{Sym}^{2}\Omega_{X}.$
@@
The second fundamental form can be constructed by applying [MR274461, Section I.3] to the exact sequence (e.g. [MR0463157, Theorem II.8.17 (2)]):
(5.4) $0\longrightarrow N^{\vee}_{X/\mathbb{P}^{n}}\longrightarrow\Omega_{\mathbb{P}^{n}}|_{X}\longrightarrow\Omega_{X}\longrightarrow 0.$
In [einniu, MP24], the $m^{\mathrm{th}}$ fundamental forms are constructed using the Taylor series map applied to the bundles of principal part $\mathcal{P}_{X/k}^{m}(\mathcal{O}_{X})$ [MR238860, Définition (16.3.6)], and using again the properties of the kernel. They suppose $k$ to be algebraically closed, but this hypothesis is not really necessary, at least in order to define the fundamental forms [^2] Confirmed by Wenbo Niu by a personal communication.. Our approach follows [MR1308020, Appendix A].
@@
In Proposition 5.3, we cannot drop the hypothesis that $X$ admits a theta characteristic. For example, a non-singular real plane curve of degree $4$ with no real points has no theta characteristics [MR286136, Section 5]. See also [MR631748, Section 5] for a thorough discussion about theta characteristics of real curves, and [MR2672429] for a characterization of the existence of theta characteristic in terms of the splitting of the stable motivic homotopy type of the curve. If $k$ is algebraically closed, then any non-singular curve admits theta characteristics [MR292836, Section 4].
@@
For arbitrary fields note that the functor $\otimes_{k}\bar{k}$ is exact, so we may suppose $k$ algebraically closed. Indeed, Serre’s intersection formula is preserved, see [MR0463157, Appendix A]. A geometric interpretation of the degenerate locus of   II as the locus of inflectional points is given by [MP24, Section 3] (see also [MR1308020, Appendix B, Theorem 2.3] and [MR506323]).
@@
Since $\mathcal{F}\to X$ is a line bundle over a curve, its Euler number is hyperbolic by [FourLines, Proposition 19]. It thus suffices to compute the rank of $e(\mathcal{F})$, which is given by the integral Euler number over an algebraically closed field. The claim now follows from the classical count of 3d(d2)3d(d-2) inflectional points to a smooth complex plane curve of degree $d$ [MR0463157, Exercise IV.2.3]. ∎
@@
The latter equation follows from the first by the Poincaré–Hopf Theorem (see e.g. [BW23, Theorem 1.1]) and the fact that $\langle 3\rangle\mathbb{H}=\mathbb{H}$, so it remains to treat the first equation. Using the same notation at the beginning of this section, assume $X=\operatorname{Spec}(A)$ is localized at a smooth $k$-point. The ideal II is free of rank one, so it is generated by $\mathrm{d}z:=z\otimes 1-1\otimes z$ for some $z\in A$. There exists an isomorphism
@@
In our case, $\partial_{z}^{m}b=D_{z}^{(m)}b$ are the Hasse derivatives of the regular element $b$ with respect to the local parameter $z$ (see [MR1308020, Note 6.4.2]). Thus, in this particular open subset, the second and third fundamental forms are given by $(\partial_{z}^{2}b)\mathrm{d}z\otimes\mathrm{d}z$ and $(\partial_{z}^{3}b)\mathrm{d}z\otimes\mathrm{d}z\otimes\mathrm{d}z$, respectively. Since $\frac{\partial}{\partial z}D^{(n)}_{z}=(n+1)D^{(n+1)}_{z}$, the Jacobian of   II on this open subset is given by
$\frac{\partial}{\partial z}((\partial^{2}_{z}b)\mathrm{d}z\otimes\mathrm{d}z)=3(\partial^{3}_{z}b)\mathrm{d}z\otimes\mathrm{d}z\otimes\mathrm{d}z.$